Q)A block of mass 100 g is moved with a speed of 5 m/s at the highest point in a closed circular tube of radius 10 cm in a vertical plane. The cross- section of the tube is such that the block just fits in it. The block makes several oscillations inside the tube and finally stops at the lowest point. Find the work done by the tube on the block during the process.

Grade:12

3 Answers

SOURAV CHAKRABARTY
14 years ago

.2 j

anandpal shekhawat
14 years ago
hi, Ei = 1/2mv^2 + mg2r = 1.45J Ef = 0 so, work done = 0-1.45 = -1.45J
Aditya Kumar
5 years ago
Apply wet we get
0.5m(0^2-5^2)=w by tube +w by gravity
                         =.                  +Mg×2r
And on solving
We get 
W by tube =-1.45